### v2

# A Computational Framework for Efficient Model Evaluation with Causal Guarantees

Abstract
In order to reduce the cost of experimental evaluation for models, we introduce a computational theory of evaluation for prediction and decision models: build evaluation model to accelerate the evaluation procedures. We prove upper bounds of generalized error and generalized causal effect error of given evaluation models. We also prove efficiency, and consistency to estimated causal effect from deployed subject to evaluation metric by prediction. To learn evaluation models, we propose a meta-learner to handle heterogeneous evaluation subjects space problem. Comparing with existed evaluation approaches, our (conditional) evaluation model reduced 24.1%-99.0% evaluation errors across 12 scenes, including individual medicine, scientific simulation, social experiment, business activity, and quantum trade. The evaluation time is reduced 3-7 order of magnitude comparing with experiments or simulations.
1 Introduction
Throughout the history of computer science, manually handing the data has been replaced by computational approach, such as computational linguistic (OpenAI et al. (2024), DeepSeek-AI et al. (2025)), computational biology (Abramson et al. (2024)), and computational learning (Valiant (1984)). Here we extend the computational theory to the ubiquitous domain of evaluation.
Randomized controlled trials (Fisher (1974), Box et al. (2005)) are widely regarded as the gold standard for evaluating the efficacy of interventions, such as drugs and therapies, due to their ability to minimize bias and establish causal relationships between interventions and outcomes. By randomly assigning participants to intervention or control groups, RCTs effectively balance both observed and unobserved confounding variables, thereby mitigating the influence of hidden common causes (Reichenbach (1999)). However, conducting RCTs to assess every potential intervention is often prohibitively expensive and, in certain cases, unfeasible. This challenge is particularly pronounced in fields like artificial intelligence, where models may possess hundreds of parameters, resulting in an infinite and high-dimensional space of potential interventions. Evaluating such models through randomized experiments is resource-intensive, demanding substantial investments of time, finances, personnel, and materials (Speich et al. (2018)). Moreover, the extended duration of these experiments hampers the rapid iteration and optimization of models, rendering the process inefficient. Given these limitations, alternative methodologies that balance rigorous evaluation with practical feasibility are essential, especially in rapidly evolving domains requiring swift assessment and deployment of new models.
An improvement is to build the centralized A/B platform, which has been proven successes in application updates and recommendation in Byte Dance (Byte-Dance ). The implementation of centralized A/B testing platforms can significantly reduce the costs associated with conducting randomized trials for researchers. However, this cost reduction does not inherently enhance the utility derived from each individual trial and may inadvertently increase the risk to participants. Notably, the lower cost and increased accessibility of such platforms can lead to a substantial rise in the number of randomized trials conducted. This proliferation of experiments raises ethical concerns regarding participant exposure to potential harm during the randomization process. For instance, as highlighted by Zhou et al. (2024b), the automation and scalability of A/B testing frameworks, while beneficial for rapid experimentation, necessitate careful consideration of ethical implications to safeguard participant well-being. Polonioli et al. (2023) emphasizes the need for ethical and responsible experimentation to protect users and society.
Evaluatology (Zhan et al. (2024)) emerges as an independent filed in recent years. However, there is still a lack of a theory to help us find an effective and efficient evaluation system that satisfies the concerns and interests of the stakeholders. Here, we introduce a computational theory to evaluate the effect of given evaluation subjects. The main benefit to evaluate by computation is its efficiency and low cost. If we learned the connection between the experiment result and pre-experiment subjects, the evaluation cost will be reduced dramatically comparing with experiments.
Figure 1: Procedure of computational evaluation. C is evaluation condition, S is evaluation subject, M is evaluation metric.
The procedure of our computational evaluation is listed in figure 6. First, data of evaluation condition, evaluation subject, and evaluation metric are collected. If causal relationship is needed, the subject is required to be independently, randomly, identically sampled (IRIS) so that the causal effect from the subject to the metric can be bounded in the following steps. Next, evaluation model is learned from the collected data, and generalized evaluation error is upper bounded under independently, identically distributed error (IIDE) assumption. Also, the evaluation model is assumed conditional unbiased to bound the causal effect from subject to metric. Then, learned evaluation model is applied to the downstream applications once the upper bounds are low enough to satisfy the real need. The mentioned assumptions are assessed by testing their necessary conditions.
In addressing on on the heterogeneous subject space problem in evaluation model learning, we propose a meta algorithm to parameterize subjects, learn (conditional) evaluation models, and inference separately. The upper bounds of generalized error for both metric and its first order difference on subject are minimized by minimizing the empirical error. Also, learning evaluation model from evaluation units directly is very challenging in some cases. So we add a proxy module for the subject in addressing the challenge. The proxy module will create some proxy metrics according some existing computational evaluation methods, which can help us to improve the performance of evaluation model.
In our evaluation experiment, we test evaluation model on 11 scenes including individual medicine, scientific simulation, insurance, advertisement, trade. Comparing with the existed computational evaluation methods, our method reduced 24.1%-99.0% evaluation errors while remain the advantage of computational evaluation efficiency. We also test our conditional evaluation model in the quantitative trade scene. Comparing with baseline backtesting strategy, our evaluation model reduced 89.4% evaluation errors. The evaluation acceleration ratio is ranged from 1000 times to 10000000 times comparing with experimental or simulation based evaluation in the 12 scenes.
2 Notation
Variable Name Description
$S$ evaluation subject the operable and indivisible unit that we
want to deploy, remove, understand, and modify
$C$ evaluation condition set of pre-evaluation features of evaluation subject,
including inner attributes and environments
that identify a real evaluation process
$M$ evaluation metric subset of stakeholders’ buy-in metrics
$EM$ or $\hat{f}$ evaluation model a function that takes an evaluation subject and
evaluation conditions as inputs and
produces evaluation metrics as its output
L evaluation error function a function to measure the discrepancy
(ranged from 0 to 1) between output of evaluation model and true metric
$E_{emp}$ empirical evaluation error average evaluation error of empirical sample
$E_{gen}$ generalized evaluation error expected evaluation error of super-population
Table 1: Notation table for used variables.
Table 1 shows the notation we used in this paper. In order to simplify the evaluation problem, evaluation metric and difference of evaluation metric are assumed as measurable and pairwise comparable. The computational evaluation problem is modeled as computing the evaluation metrics for evaluation subjects with certain evaluation conditions.
3 Theoretical analysis
3.1 Upper bound of generalized evaluation error
Theorem 1.
Upper bound. Given any evaluation model $\hat{f}$, $P(E_{gen}(\hat{f})<E_{emp}(\hat{f})+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})})\geq 1-\sigma$ where $n$ is number of independent identical distributed error (IIDE) measurements, $10<1-\sigma<$ 1 is confidence.
In Theorem 1, we show that the generalization error $E_{gen}$ is bounded by the empirical error $E_{emp}$. the number of error measurements $n$, and the significance level $1-\sigma$ under the IIDE assumption. This result underpins our ability to use a finite number of evaluation samples to compare the performance of evaluation models across an infinite subject and condition space. For example, when the evaluation subject space is extremely large or infinite—as in the cases of AI diagnosis models, AI treatment models, or quantitative strategy models—Theorem 1 allows us to generalize our empirical analysis to the entire subject space. Similarly, in evaluation condition spaces, such as those involving AI-driven discovery of unseen materials based on molecular and structural conditions, the generalization errors for any evaluation model can be bounded using the same theorem.
Once the upper bound is smaller than given error tolerance, evaluation model $f$ is regarded to meet the practical needs with high probability $1-\sigma$. Additionally, Theorem 1 provides a means to estimate the number of error measurements (or evaluation samples) required prior to sampling.
We emphasize that IIDE can also be satisfied if the non-IID components of $(C,S,M)$ was offset by the evaluation model or the evaluation error function although independent identical distributed $(C,S,M)$ is a sufficient condition of IIDE. In fact, the assumption of independent and identically distributed samples is considered excessively strong in this paper, particularly with respect to IID conditions and IID outcomes.
3.2 Hidden common cause
The hidden common cause between subject and metric may introduce the confounding bias if the evaluation task is causality sensitive, such as individual medicine. Causal inference is often used to handle hidden common cause from data. However, applying existed methods to handle hidden common cause problem in evaluation faced those challenges:
  • Hidden identical common cause (HIC). If the assignment of $S$ is determined by an unlisted identical common cause $S:=E$, and there exists effect from $E$ to $EI$. We may mistakenly attributing the effect on $EI$ of $E$ to $S$. The identical hidden confounder can never be excluded from statical analysis due to the identical property if $S$ was never controlled by randomization. Therefore, effect from $S$ to $EI$ is unidentifiable in all kinds of non-intervention based analysis, including diagram-based identification (Pearl (1995), Tian and Pearl (2002), Shpitser and Pearl (2006), Yan (2025)), observation study (ROSENBAUM and RUBIN (1983)), causal representation learning (Schölkopf et al. (2021)), and causal discovery (Spirtes and Glymour (1991), Zhang and Hyvärinen (2009)).
  • High-dimensional evaluation subject. Potential outcomes (Splawa-Neyman et al. (1990), Holland (1986), Imbens and Rubin (2015)) often assumes each evaluation subject must be sampled once which often do not established for high-dimensional evaluation subjects. For instance, parameters of AI model can be hundreds or thousands and more complex which structure is different from continuous dose-level analysis with low dimension.
For causality sensitive task, any hidden common cause between subject and metric is excluded including HIC if the subject was randomly, independently, identically sampled or assigned as shown in theorem 2, 3, and 4. For high-dimensional evaluation subject, the causal effect of evaluation error is bounded and evaluation result can be generalized to the whole subject space whatever how large the space size as shown in theorem 4.
Theorem 2.
Strict causal advantage. Given $\{c_{i},s_{i},m_{i}\}_{i=1}^{n}$, where $c$ is evaluation condition, $s$ is independent from $c$, and $e$, and $s$ is independently, randomly, identically sampled from distribution $Pr(S)$, $m$ is generate by a unknown and inaccessible function $m=f(c,e,s)$, $e$ is other unlisted random variable, evaluation error function $L$ is root mean square error, $\forall f_{1}$, $\forall f_{2}$ where $f_{1}$ and $f_{2}$ are unbiased estimation of $f$ given $c$, we have Efficiency: if $E_{gen}(f_{1})<E_{gen}(f_{2})$ for any $s$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $E_{gen}(f_{1}(c,s_{a})-f_{1}(c,s_{b}))<E_{gen}(f_{2}(c,s_{a})-f_{2}(c,s_{b}))$; Consistency: if $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f})=0$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f}(c,s_{a})-\hat{f}(c,s_{b}))=0$.
From theorem 2, minimization of generalized mean square causal effect evaluation error can be performed by minimizing generalized evaluation error greedily until the evaluation error is close to 0.
Theorem 3.
Causal bound with positivity. $\forall s_{a}\in S,\forall s_{b}\in S$, given any unbiased evaluation model $\hat{f}(c,s_{a})$ and $\hat{f}(c,s_{b})$,
$P(E_{gen}(\Delta\hat{f})<2\max\{E_{emp}(a)+\sqrt{\frac{1}{2n_{a}}\ln(\frac{2}{\sigma}})),E_{emp}(b)+\sqrt{\frac{1}{2n_{b}}\ln(\frac{2}{\sigma})}\})\geq 1-\sigma$
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n_{a}$ and $n_{b}$ are number of independently, identically distributed error (IIDE) measurements where $s=s_{a}$ and $s=s_{b}$, $10<1-\sigma<$ 1 is confidence, $s$ is independently, randomly, identically sampled (IRIS), evaluation error function $L$ is mean square error ranged from 0 to 1.
Theorem 4.
Causal bound with non-positivity. Given unbiased evaluation model $\hat{f}$, then $\forall s_{a}\in S,\forall s_{b}\in S$,
$P(E_{gen}(\Delta\hat{f})<2(E_{emp}(\hat{f})+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})}))\geq 1-\sigma$
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n$ is number of independently, randomly, identically sampled error measurements (IIDE), $10<1-\sigma<$ 1 is confidence, $s$ is independently, randomly, identically sampled (IRIS), evaluation error function $L$ is mean square error ranged from 0 to 1.
Theorem 3 provides the rationale for why our evaluation models can learn the causal effect from subjects to metrics rather than mere correlations when the evaluation subjects are independently, randomly, and identically sampled (IRIS), the evaluation models are unbiased, and the evaluation errors for are independently, randomly, identically distributed (IIDE). In practice, $n_{a}$ and $n_{b}$ are typically zero for infinite evaluation subject (non-positivity Imbens and Rubin (2015)); in such cases, upper bound on the causal effect error is given by Theorem 4. Proofs of all the theorems and specific upper bounds were given in appendix A, G.
Comparing with Saito and Yasui (2020) and Alaa and Van Der Schaar (2019), our unbiased assumption has potential to be rejected from real data testing as shown in the experiment part while they did not provide approaches to reject their unbiased causal effect assumption (Saito and Yasui (2020)) or min-max optimal effect estimator assumption (Alaa and Van Der Schaar (2019)) from real data without counterfactual. Also, IID samples assumption in their work is too strong for causal evaluation, especially IID conditions and IID outcomes.
4 Learning evaluation models
In evaluation processes, two critical factors are evaluation error and evaluation cost. Our objective is to reduce evaluation cost while maintaining an acceptable level of generalized evaluation (causal) error through computational evaluation. To achieve this, we can minimize empirical evaluation error because upper bound size is fixed given number of error measurements $n$ and confidence requirement $1-\sigma$ in theorem 1.
Directly learning an evaluation model from evaluation samples presents significant challenges. To address this, we introduce a proxy module that generates surrogate metrics using established computational methods, such as dataset-based approaches including bootstrapping, hold-out validation, and cross-validation. These techniques enhance the performance of the evaluation model by providing more robust and reliable features. Furthermore, to effectively manage heterogeneous subjects, we employ vectorization strategies, develop specialized sub-models, and conduct inferences tailored to each subject type. This approach ensures that the evaluation model can accommodate the diverse characteristics inherent in heterogeneous subject spaces.
(a) Evaluation model learning.
(b) Evaluation model inference.
(c) Evaluation model Selection.
Figure 2: Learn evaluation models from data.
The proposed meta-learning algorithm for evaluation model is illustrated in Figure 2. Initially, data are classified based on subject type. Subsequently, distinct vectorization methods are applied corresponding to each subject type. Proxy metrics are then computed using predefined proxy functions, facilitating the development of tailored evaluation models. For the assessment of unseen subjects, the process involves computing proxy values, determining the subject type, selecting the appropriate vectorization function, and inputting these into the corresponding evaluation model to predict relevant metrics. The base learners for the evaluation models can vary and include algorithms such as linear models, Random Forests (RF), Multi-Layer Perceptrons (MLP), CatBoost, XGBoost, and LightGBM. Following the development of these evaluation models, an upper bound is calculated to check whether the models meet practical requirements and constraints.
5 Assessing assumptions
The IRIS assumption can be satisfied in data collection while the other assumptions may not be satisfied in real world. In order to assess the IIDE and unbiased assumptions, statistical analyses can be employed to verify necessary conditions under those assumptions, providing credibility to the theorems derived from this assumption for someone. It is important to note that these assessments do not conclusively prove the unbiasedness assumption. Instead, they serve as methods to potentially refute the assumption under specific conditions, rather than confirming its validity.
5.1 Assessing IIDE assumption
In highly controlled environments, such as CPU evaluation, the assumption can be rigorously satisfied by independently and randomly selecting tested chips and configurations from an identical distribution. However, in many open environments, the IIDE assumption is not directly testable on real-world data, such as individual medicine, and social experiment.
First, the normalized means of error subsets should follow a common Gaussian distribution, as implied by the central limit theorem. To test this, we randomly generate 30 subsets and apply the D’Agostino-Pearson test to assess whether their normalized means adhere to a single Gaussian distribution. Second, error distributions across different splits should be identical. We randomly partition the errors into 30 subsets and use the Kolmogorov-Smirnov test to determine whether the distributions across these subsets are statistically indistinguishable.
5.2 Assessing Unbiased assumption
Our bounds requires unbiased evaluation models while sometime the evaluation model may not be unbiased. To assess the unbiasedness of the evaluation model, we employ the following statistical procedures: 1) Global Unbiasedness Assessment: We conduct a one-sample t-test on the entire dataset to determine if the model’s errors have a mean of zero, indicating unbiasedness across the universal set; 2) Subset Unbiasedness Assessment: We randomly select 30 subsets from the dataset and perform one-sample t-tests on each to evaluate whether the mean errors within these subsets deviate significantly from zero. This step examines the consistency of the model’s unbiasedness across different data partitions.
6 Experiment
We performed experiments for evaluation model without condition and evaluation model with condition respectively in 12 scenes of model evaluation, including randomized experiment prediction, scientific simulation, advertising, insurance, and quantum trade. The detailed evaluation scenarios (including experiment setting), assumption assessing, hyperparameter tuning, and evaluation results are listed in the appendix B, C, D, E.
6.1 Evaluation model
To demonstrate the generalizability of our evaluation model in subject space, we consider 11 distinct scenarios. In these scenes, computational evaluation can be helpful to reduce experimental cost, shorten simulation time, or enhance the efficiency of advertising and insurance sales.
Although subjects are independently, randomly, and identically sampled from the subject space, they cannot be randomly deployed in real-world settings for direct measurement of benefit-related metrics (e.g., mortality, rehabilitation time). Instead, our evaluation system defines the evaluation metric as the post-experiment metric on unseen test data with randomized treatment. This metric captures the causal effect from subject to post-experiment outcome, as established in Theorem 4. Empirical studies (Ai et al. (2022); Zhou et al. (2023, 2024a); He et al. (2024); Gentzel et al. (2021)) have demonstrated a strong correlation between the decision model’s predictive performance on randomized experimental data and its performance following randomized deployment. Accordingly, we employ these predictive metrics to approximate randomized post-deployment performance, which can then be integrated into the decision via an appropriate utility function.
(a) Normalized negative RMSE when evaluation metric is ROC-AUC.
(b) Normalized negative RMSE when evaluation metric is ACC.
(c) Normalized negative RMSE when evaluation metric is RMSE.
(d) Normalized negative RMSE when evaluation metric is $R^{2}$.
Figure 3: Normalized empirical negative RMSE of evaluation models crossing scenes (6 scenes in figure 3(a) and figure 3(b), 5 scenes in figure 3(c) and figure 3(d). The confidence level of confidence interval bar is set as 95%. The baselines are holdout 100%, holdout 50%, holdout 10%, 5-fold cross-validation, 10-fold cross-validation, and bootstrap. We test different base learners (Linear, MLP, SVM/SVR, Random forest, XGBoost, LighGBM, and CatBoost) for heterogeneous subject space.
Across the 11 scenarios, we compare our evaluation model against six existing validation methods that do not incorporate prior information from data. The normalized negative empirical RMSE is illustrated in Figure 3. Our best evaluation model achieves error reductions of 94.7–99.0% in RMSE, 81.7–95.5% in $R^{2}$, 24.1–77.0% in ROC-AUC, and 39.2–89.5% in ACC relative to the holdout-100 validation method. This performance improvement is attributable to the effective modeling of the relationships among the subject, its proxy metrics, and the evaluation metrics. Notably, models that capture spurious correlations (MLP and SVR without hyperparameter tuning) fail to reduce evaluation error.
6.2 Conditional evaluation model
In order to demonstrate the generalizbility of conditional evaluation model on the condition space further, we introduce the trade backtesting task of A-share market in China to test the performance of conditional evaluation model. In this scene, the computational evaluation can be helpful to exclude the influence of hidden common cause between trade strategy and return of invest (RoI), and to reduce the evaluation cost of trade strategy’s deployment.
The evaluation subjects are trade models whose inputs are last day’s variables of the stock and output are one of three decisions (buy 1 hand, sell 1 hand, and hold) at open this day. The metric is Return of Invest of a subject in a time slot (10 days) in future. The condition is the stock’s time-series variables in the last time slot. In this scene, we not only measure the subject and its metric, but also consider the pre-trade variables (condition) of the subject in an attempt to reduce evaluation error further. We use the unseen future data and performance of unseen subjects for conditional evaluation model testing.
Similarly, we build the evaluation systems by combining the real data and a float model whose input is subject vector and last day’s variables of the stock, and output is a small float ratio of opening price. We want to use it to simulate the exclusion of hidden common cause on the stock market by the data without the randomized deployment of the subject.
Backtesting method RMSE of RoI
Baseline (Last10Days) 3.067
HetEM (Linear) 2.163
HetEM (CatBoost) 0.527
Table 2: Performance of backtesting methods.
We compare our conditional evaluation models with a baseline backtesting methods as shown in the table 2. The reason we did not take the time cross-validation, k-fold cross-validation, and combinatorial symmetry cross-validation (Bailey et al. (2013)) into baselines is that those methods are train-valid fused rather than evaluation-targeted methods. When the validation part is separated, they will give the same trivial constant prediction which can not utilize the evaluation conditions for each time slot.
Our best conditional evaluation models (HetEM-CatBoost) for heterogeneous subjects reduced 89.4% evaluation errors comparing with the baseline method. Even the linear conditional evaluation model can reduce the 34.0% RMSE of estimated RoI. The average RMSE of estimated RoI in a future time slot is about 0.527, which may be reduced further when adding more factors of stocks. From the result of different evaluation methods, the computational evaluation approach has higher potential to solve the backtest over-fitting problem (Bailey et al. (2013)) in stock market.
6.3 Interpretability
(a) Shapley Value of Subject and Proxy when metric is ROC-AUC.
(b) Shapley Value of Subject and Proxy when metric is ACC.
(c) Shapley Value of Subject and Proxy when metric is RMSE.
(d) Shapley Value of Subject and Proxy when metric is $R^{2}$.
Figure 4: Shapley value of subject vector and proxy metrics on 11 scenes (6 scenes in figure 4(a) and figure 4(b), 5 scenes in figure 4(c) and figure 4(d). The outcome of null set is set as the performance of holdout-100, and other outcomes are the negative RMSE of the linear evaluation models.
In order to understand the contribution of subject vector, and proxy metric in the evaluation of our best evaluation model respectively, we visualize the Shapley value (Shapley (1997), Lundberg and Lee (2017)) for subject vector and proxy metric of all models as shown in figrue 4. The null set’s value is set as the negative RMSE of holdout-100. From the visualization, the main contribution for heterogeneous subject whose metrics are ROC-AUC and ACC, is from the relationship between the proxy metric and true metric despite of an example in climate prediction scene. For heterogeneous subjects whose metrics are RMSE and $R^{2}$, the contribution of subject and proxy metric is almost equal to each other. It reveals the effectiveness of the introduced proxy module and subject vectorization module in our learning algorithm respectively.
In order to understand the contribution of the condition vector, subject vector, and proxy metrics in the evaluation of our best conditional evaluation model, we use Shapley values to explain the contribution of modules (Shapley (1997), Lundberg and Lee (2017)). The null set is set as the baseline backtesting method. The contribution rate of condition vector, subject vector, and proxy metrics for the HetEM(CatBoost) are 28.8% (0.729 uplift RMSE of RoI), 35.6% (0.905 uplift RMSE of RoI), and 35.6% (0.906 uplift RMSE of RoI) respectively. It reveals the importance of evaluation condition to reduce the evaluation errors in the A-share trade scenes of China.
6.4 Evaluation cost
The primary advantage of our computational evaluation method is its efficiency and low cost. Rather than conducting an unbounded number of experiments for infinitely many subjects (e.g., AI models or other agents), our approach only requires the collection of initial data, training an evaluation model, and subsequently applying it in real-world scenarios.
The acceleration ratio of time achieved by our computational method ranges from 1,000 to 10,000,000 per evaluation subject in the 12 scenarios as shown in appendix. Although experimental and simulation-based evaluations can be accelerated via parallel processing (excluding additional evaluation costs), the same parallelization strategies can be readily applied to computational evaluation approach. The detailed comparison of evaluation time can be seen in the appendix F.
7 Limitation and conclusion
Limitation. Despite the theoretical contributions and empirical results, the proposed approach is subject to several limitations. First, our theoretical guarantees rely on the IIDE, IRIS, and unbiasness assumptions, which may not hold in all real-world scenarios although statistical tests provide some support. Second, while our method has been validated across multiple application domains, its scalability and generalization in extremely high-dimensional settings require further investigation. Third, lower error and tighter bound are required to apply evaluation models into real applications.
Conclusion. In this work, we propose a novel computational framework for the evaluation of decision models that rigorously derives upper bounds on generalized evaluation error and causal effect error, and employs a meta-learning strategy to address heterogeneity in evaluation subject spaces. Our extensive experimental results, conducted over 12 diverse scenarios—including individual medicine, scientific simulation, social experiments, business activity, and quantum trade—demonstrate significant error reductions (ranging from 24.1% to 99.0%) and substantial acceleration (up to $10^{7}$) compared to traditional evaluation methods. These findings highlight the potential of our technical route to drastically reduce evaluation costs while enabling rapid model iteration. Future work will focus on scaling the framework to larger domains and enhancing the validation of key assumptions to further improve robustness and applicability. The paradigm of computational evaluation can be seen in the appendix H.
Appendix A Proofs
Theorem 1.
Upper bound. Given any evaluation model $\hat{f}$, $P(E_{gen}(\hat{f})<E_{emp}(\hat{f})+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})})\geq 1-\sigma$ where $n$ is number of independent identical distributed error (IIDE) measurements, $10<1-\sigma<$ 1 is confidence.
Proof.
Denote $\epsilon_{0}\cong\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})}$, inequality in theorem 1 is equivalent to
$P(E_{gen}(\hat{f})-E_{emp}(\hat{f})<\epsilon_{0})\geq 1-\sigma$
. Because event $E_{gen}(\hat{f})-E_{emp}(\hat{f})<\epsilon_{0}$ and event $E_{gen}(\hat{f})-E_{emp}(\hat{f})\geq\epsilon_{0}$ are complementary events, so
$P(E_{gen}(\hat{f})-E_{emp}(\hat{f})<\epsilon_{0})=1-P(E_{gen}(\hat{f})-E_{emp}(\hat{f})\geq\epsilon_{0})$
. Bring it in, we have
$1-P(E_{gen}(\hat{f})-E_{emp}(\hat{f})\geq\epsilon_{0})\geq 1-\sigma$
. Then it can be simplified as
$P(E_{gen}(\hat{f})-E_{emp}(\hat{f})\geq\epsilon_{0})\leq\sigma$
. Because evaluation error is distributed on $[0,1]$, and Hoeffding inequality Hoeffding [1963],
$P(E_{gen}(\hat{f})-E_{emp}(\hat{f})\geq\epsilon_{0})$
$\leq 2e^{-\frac{2\epsilon_{0}^{2}n^{2}}{\sum_{i=1}^{n}(1-0)^{2}}}=2e^{-\frac{2(\sqrt{\frac{1}{2n}(\ln(\frac{1}{\sigma}))})^{2}n^{2}}{\sum_{i=1}^{n}(1-0)^{2}}}=\sigma$
Theorem 2.
Strict causal advantage. Given $\{c_{i},s_{i},m_{i}\}_{i=1}^{n}$, where $c$ is evaluation condition, $s$ is independent from $c$, and $e$, and $s$ is independently, randomly, identically sampled from distribution $Pr(S)$, $m$ is generate by a unknown and inaccessible function $m=f(c,e,s)$, $e$ is other unlisted random variable, evaluation error function $L$ is root mean square error, $\forall f_{1}$, $\forall f_{2}$ where $f_{1}$ and $f_{2}$ are unbiased estimation of $f$ given $c$, we have Efficiency: if $E_{gen}(f_{1})<E_{gen}(f_{2})$ for any $s$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $E_{gen}(f_{1}(c,s_{a})-f_{1}(c,s_{b}))<E_{gen}(f_{2}(c,s_{a})-f_{2}(c,s_{b}))$; Consistency: if $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f})=0$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f}(c,s_{a})-\hat{f}(c,s_{b}))=0$.
Proof.
$\forall k=1,2$, denote $\epsilon_{k}(c,s,e)=f_{k}(c,s)-f(c,e,s)$, according to conditional unbiasness, we have $E(\epsilon_{k}(c,s))=0$. The error is denoted as
$\Delta f_{k}(s)=f_{k}(c,s_{a})-f_{k}(c,s_{b})-f(c,s_{a},e)-f(c,s_{b},e)=\epsilon_{k}(c,s_{a},e)-\epsilon_{k}(c,s_{b},e)$
, the square error is denoted as
$(\Delta f_{k})^{2}=(\epsilon_{k}(c,s_{a},e)-\epsilon_{k}(c,s_{b},e))^{2}$
, the mean square error is denoted as
$E((\Delta f_{k})^{2})=E(\epsilon_{k}(c,s_{a},e)^{2})+E(\epsilon_{k}(c,s_{b},e)^{2})-2E(\epsilon_{k}(c,s_{a},e)\epsilon_{k}(c,s_{b},e))$
. According to the condition unbiasness and IRIS assumption, the third term
$E(\epsilon_{k}(c,s_{a})\epsilon_{k}(c,s_{b}))=E_{c}(E(\epsilon_{k}(c,s_{a},e)\epsilon_{k}(c,s_{b},e)|c))$
$=E_{c}(E(\epsilon(c,s_{a})|c)E(\epsilon(c,s_{b})|c))=E_{c}(0*0)=E_{c}(0)=0$
, then because $\forall s$, $E_{gen}(f_{1})<E_{gen}(f_{2})$, that is $E(\epsilon_{1}(c,s,e)^{2})<E(\epsilon_{2}(c,s,e)^{2})$ for any $s$, then
$E((\Delta f_{1})^{2})=E(\epsilon_{1}(c,s_{a},e)^{2})+E(\epsilon_{1}(c,s_{b},e)^{2})<E(\epsilon_{2}(c,s_{a},e)^{2})+E(\epsilon_{2}(c,s_{b},e)^{2})=E((\Delta f_{2})^{2})$
, which can be write as $E_{gen}(f_{1}(c,s_{a})-f_{1}(c,s_{b}))<E_{gen}(f_{2}(c,s_{a})-f_{2}(c,s_{b}))$ when error function is mean square error. Following, the consistency can be proved spontaneously by the Squeeze Theorem in Calculus. ∎
Theorem 3.
Causal bound with positivity. $\forall s_{a}\in S,\forall s_{b}\in S$, given any unbiased evaluation model $\hat{f}(c,s_{a})$ and $\hat{f}(c,s_{b})$,
$P(E_{gen}(\Delta\hat{f})<2\max\{E_{emp}(a)+\sqrt{\frac{1}{2n_{a}}\ln(\frac{2}{\sigma}})),E_{emp}(b)+\sqrt{\frac{1}{2n_{b}}\ln(\frac{2}{\sigma})}\})\geq 1-\sigma$
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n_{a}$ and $n_{b}$ are number of independently, randomly, identically sampled error measurements where $s=s_{a}$ and $s=s_{b}$, $10<1-\sigma<$ 1 is confidence, $s$ is independently, randomly, identically sampled (IRIS), evaluation error function $L$ is mean square error.
Proof.
Because evaluation error function is mean square error,
$E_{gen}(\Delta\hat{f})\cong E((\Delta\hat{f}-\Delta f)^{2})=E((\hat{f}(c,s_{a})-\hat{f}(c,s_{b})-(f(c,s_{a},e)-f(c,s_{b},e)))^{2})$
. Denote $\epsilon_{a}(c,e)\cong\hat{f}(c,s_{a})-f(c,s_{a},e)$ and $\epsilon_{b}(c,e)\cong\hat{f}(c,s_{b})-f(c,s_{b},e)$,
$E_{gen}(\Delta\hat{f})=E(\epsilon_{a}(c,e)^{2}+\epsilon_{b}(c,e)^{2}-2\epsilon_{a}(c,e)\epsilon_{b}(c,e))$
. Due to properties of expectation,
$E_{gen}(\Delta\hat{f})=E(\epsilon_{a}(c,e)^{2})+E(\epsilon_{b}(c,e)^{2})-2E(\epsilon_{a}(c,e)\epsilon_{b}(c,e))$
. Due to IRIS assumption, the sub population of errors where $s=a$ and $s=b$ is also independently sampled,
$E_{gen}(\Delta\hat{f})=E(\epsilon_{a}(c,e)^{2})+E(\epsilon_{b}(c,e)^{2})-2E(\epsilon_{a}(c,e))E(\epsilon_{b}(c,e))$
. Due to the unbiased assumption and IIDE assumption for populations where $s=s_{a}$ and $s=s_{b}$,
$E_{gen}(\Delta\hat{f})=E(\epsilon_{a}(c,e)^{2})+E(\epsilon_{b}(c,e)^{2})$
. Denote $\epsilon_{0}=2\max\{\nu_{a},\nu_{b}\}$ where $\nu_{a}=E_{emp}(a)+\sqrt{\frac{1}{2n_{a}}\ln(\frac{2}{\sigma}})$ and $\nu_{b}=E_{emp}(b)+\sqrt{\frac{1}{2n_{b}}\ln(\frac{2}{\sigma})}$, due to Bonferroni inequality,
$P(E_{gen}(\Delta\hat{f})<\epsilon_{0})=P(E(\epsilon_{a}(c,e)^{2})+E(\epsilon_{b}(c,e)^{2})<\epsilon_{0})$
$\geq P(E(\epsilon_{a}(c,e)^{2})<\frac{\epsilon_{0}}{2}\bigcap E(\epsilon_{b}(c,e)^{2})<\frac{\epsilon_{0}}{2})$
$\geq P(E(\epsilon_{a}(c,e)^{2})<\frac{\epsilon_{0}}{2})+P(E(\epsilon_{b}(c,e)^{2})<\frac{\epsilon_{0}}{2})-1$
$\geq P(E(\epsilon_{a}(c,e)^{2})<\nu_{a})+P(E(\epsilon_{b}(c,e)^{2})<\nu_{b})-1$
$\geq 1-\frac{\sigma}{2}+1-\frac{\sigma}{2}-1$
$=1-\sigma$
Theorem 4.
Causal bound with non-positivity. Given arbitrary conditional unbiased (CU) evaluation model $f$ given any $c$, then $\forall s_{a}\in S,\forall s_{b}\in S$,
$P(E_{gen}(\Delta\hat{f})\leq 2(E_{emp}(\hat{f})+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})}))\geq 1-\sigma$
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n$ is number of independently, randomly, identically sampled error measurements (IIDE), $10<1-\sigma<$ 1 is confidence, $s$ is independently, randomly, identically sampled (IRIS), evaluation error function $L$ is mean square error ranged from 0 to 1.
Proof.
According to same proof and denotation in theorem 3,
$E_{gen}(\Delta\hat{f})=E(\epsilon_{a}(c,e)^{2})+E(\epsilon_{b}(c,e)^{2})-2E(\epsilon_{a}(c,e))E(\epsilon_{b}(c,e))$
. Due to IRIS assumption, the sub population of errors where $s=a$ and $s=b$ is also independently, randomly and identically sampled following the whole population,
$E_{gen}(\Delta\hat{f})=2E(\epsilon(c,e)^{2})-2E(\epsilon(c,e))^{2}$
. Due to the unbiased assumption,
$E_{gen}(\Delta\hat{f})=2E(\epsilon(c,e)^{2})$
. According to theorem 1, let evaluation error function be MSE,
$P(E(\epsilon(c,e)^{2})\leq\sum_{i=1}^{n}(\hat{f}(c_{i},s_{i})-m_{i})^{2}+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})})\geq 1-\sigma$
where $m_{i}$ is true metric. Due to the property of equivalent events,
$P(2E(\epsilon(c,e)^{2})\leq 2(\sum_{i=1}^{n}(\hat{f}(c_{i},s_{i})-m_{i})^{2}+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})}))\geq 1-\sigma$
, which can be rewrite as
$P(E_{gen}(\Delta\hat{f})\leq 2(E_{emp}(\Delta\hat{f})+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})}))\geq 1-\sigma$
Appendix B Evaluation scenarios
We introduce 12 scenarios to test the performance of our (conditional) evaluation model. The evaluation subject and evaluation metric in those applied scenarios is listed in the table 3.
The heterogeneous evaluation subject space is configured as a combination of a linear subject space (logistic regression or linear regression) and a non-linear subject space (MLP Classifier or MLP Regressor). The model’s input dimension is the number of features, and output dimension is the number of the targets. The MLP’s hidden layer size limit is set as 8.
The IRIS sampling of the subject is two steps. The first step is the subject type sampling with 0.5 probability to choose linear model and 0.5 probability to choose MLP model. The second step is to choose a specific model in the sub space. For both linear model and MLP, we use Normal distribution $N(0,1)$ to determine the parameter of the models. The outcome when we calculate Shapley value is set as the negative evaluation errors.
Scene Evaluation Subject (Model) Application Metric
Individual input is patients’ record, intelligent Alert, ROC-AUC
Treatment and random pop-up, and Withdrawal ACC
output is mortality of AKI
input is patients’ record, Covid-19 recovery RMSE
and random therapy, (TERECO) $R^{2}$
output is walk distance
Scientific input is setting of fission reactor, nuclear power plant RMSE
Simulation output is k-inf and PPPF (Nuscale) $R^{2}$
input is particle features higgs detection ROC-AUC
output is whether it is higgs (Higgs) ACC
input is initial parameter of climate, climate simulation ROC-AUC
output is whether the climate crash (Climate) ACC
input is joint moving of robotic arm, robot action RMSE
output is distance to target (kin8nm) $R^{2}$
input is parameter of electric grid, grid simulation ROC-AUC
output is whether the grid crash (Grid) ACC
Social input is students’ record, individual teaching RMSE
Experiment and random class (Calculus) $R^{2}$
output is students’ grading
Business input is customers’ record, sell insurance ROC-AUC
and random recommendation (Insurance) ACC
output is buy-in decision of custom
input is feature of a group, advertisement RMSE
and random recommendation (Ad) $R^{2}$
output is buy-in decision of custom
Quantum input is stock’s feature last day, A-share trade RoI
Trade output is decision for the stock
Table 3: Detailed setting of scene.
The samples number of scenarios Alert, Withdrawal, Higgs, Climate, Grid, Insurance, TERECO, Nuscale, kin8nm, Calculus, Ad, and A-share trade are listed in table 4. All the computation is on the one 13-inch MacBook Pro 2020 with Apple M1 chip and 16GB memory.
Scenarios Train Samples Test Samples
Alert $1600*5911*0.2=1891520$ $400*5911*0.2=472880$
Withdrawal $1600*4998*0.2=1599360$ $400*4998*0.2=399840$
Higgs $1600*2000*0.2=640000$ $400*2000*0.2=160000$
Climate $1600*540*0.2=172800$ $400*540*0.2=43200$
Grid $1600*10000*0.2=3200000$ $400*10000*0.2=800000$
Insurance $1600*45211*0.2=14467520$ $400*45211*0.2=3616880$
TERECO $1600*104*0.2=33280$ $400*104*0.2=8320$
Nuscale $1600*360=576000$ $400*360$
kin8nm $1600*8192*0.2=2621440$ $400*8192*0.2=655360$
Calculus $1600*672*0.2=215040$ $400*672*0.2=53760$
Ad $1600*1143*0.2=365760$ $400*1143*0.2=91440$
Trade $160*200*30=960000$ $40*200*20=1600000$
Table 4: Evaluation sample numbers in different scenarios.
B.1 Scenarios to test evaluation model
For 11 scenes to test evaluation model, 2000 subjects was randomly sampling from heterogeneous subject space. 20% data was used to build the real systems to generate true evaluation metrics and other 80% data was used to get the proxy metrics of a subject. When the evaluation samples are collected from the built real systems, 20% samples are used for evaluation model testing, and other 80% samples are for evaluation model learning. The experiment is performed 30 times. All the null value was removed from the raw data in preprocessing.
The proxy metrics we used for Alert, Withdrawal, Higgs, Climate, Grid, Insurance are ROC-AUC, accuracy, recall, precision, $F_{1}$ score, and PR-AUC. The proxy metrics we used for TERECO, Nuscale, kin8nm, Calculus, Ad are RMSE, $R^{2}$, MAE, MAPE, and MSE.
B.2 Scenarios to test conditional evaluation model
We collected the A-share data of China from 2024-8 to 2024-11 (89 days) by an open source tool AkShare. The collected day-wise features of stocks are “Profit Ratio”, “Average Cost”, “90 Cost Low”, “90 Cost High”, “90 Episode Medium”, “70 Cost Low”, “70 Cost High”, “70 Episode Medium”, “Closing”, “Opening”, “Highest”, “Lowest”, “Transaction Volume”, “Transaction Amount”, “Amplitude”. The strategy is to buy 1 hand, sell 1 hand, or hold by the subject following the open price of the market. The slippage and commission is not considered in this work. If the stock was limit up (limit down), then we can only hold even the model decision is to buy (sell). The subject’s input is the feature of the stock last day and output is the decision this day. The subject will be deployed and run for 1 time slot (10 trading days, about 2 weeks), and the evaluation condition is feature of the stock in the last time slot. At the end of deployment in a time slot, we will sell all the hold stocks. When the end of the slot is limit down at open, we assume that it will be sold by 90 percents of open price in future.
In order to simulate the direct effect from the subject itself to the sensitive market, we add a float model which input is the vectorized subjects and output is float rate of open price ranged from -0.1% to +0.1%.
We use data of 200 stocks with minimal ID from all 3214 stocks due to our computational limit. The start of train day is from day 10 to day 41, the start of test day is from 41 to 61. The sampling number of subject is 200. The experiment was performed 5 times to reduce the influence of randomness. The proxy metric is the last 10 days’ RoI.
Appendix C Assumption assessment
Table LABEL:tab:assessregression and table LABEL:tab:assessclass are assumption assessment results of 7 kinds of decision models for continuous and discrete target respectively. From the tables, we can see that the IIDE assumption (IID check and ID check) and unbiased evaluation model assumption (Bias Check and GroupBias Check) was not violated obviously with a high probability (>0.8) except the cases that using MLP and SVR as base regression learners of HetEM.
Table 5: The ratio that the p-value of test is larger than or equal to 0.05 in the 30 repeated experiments for regression base learners. The name of each row is “BaseLearnerName”-“TestName”-“MetricName”. The name of each column is the scene name.
Ratio $p>=0.05$ TERECO Calculus kin8nm nuscale ad
Linear-IID-RMSE 0.93 0.87 0.97 0.9 0.97
Linear-IID-$R^{2}$ 0.93 1.0 0.9 0.93 0.83
Linear-ID-RMSE 1.0 0.47 0.87 1.0 0.93
Linear-ID-$R^{2}$ 0.97 0.43 0.93 0.97 0.87
Linear-Bias-RMSE 0.9 0.43 0.77 0.8 0.77
Linear-Bias-$R^{2}$ 0.9 0.33 0.8 0.77 0.77
Linear-GroupBias-RMSE 0.92 0.85 0.87 0.87 0.85
Linear-GroupBias-$R^{2}$ 0.90 0.85 0.86 0.84 0.89
MLP-IID-RMSE 1.0 0.93 0.83 1.0 0.9
MLP-IID-$R^{2}$ 0.9 0.97 0.93 0.97 0.97
MLP-ID-RMSE 0.73 0.37 0.87 0.73 0.77
MLP-ID-$R^{2}$ 0.9 0.3 0.77 0.77 0.67
MLP-Bias-RMSE 0.5 0.26 0.57 0.57 0.5
MLP-Bias-$R^{2}$ 0.57 0.17 0.53 0.57 0.57
MLP-GroupBias-RMSE 0.68 0.40 0.86 0.70 0.72
MLP-GroupBias-$R^{2}$ 0.70 0.31 0.71 0.67 0.68
SVR-IID-RMSE 0.97 0.97 0.97 1.0 0.87
SVR-IID-$R^{2}$ 0.93 0.9 1.0 0.9 0.97
SVR-ID-RMSE 0.0 0.0 0.27 0.4 0.1
SVR-ID-$R^{2}$ 0.0 0.0 0.0 0.2 0.0
SVR-Bias-RMSE 0.0 0.0 0.1 0.17 0.03
SVR-Bias-$R^{2}$ 0.0 0.23 0.0 0.03 0.0
SVR-GroupBias-RMSE 0.0 0.01 0.19 0.34 0.11
SVR-GroupBias-$R^{2}$ 0.0 0.44 0.00 0.19 0.02
RF-IID-RMSE 0.97 0.93 0.97 0.97 0.93
RF-IID-$R^{2}$ 1.0 0.97 0.97 1.0 0.97
RF-ID-RMSE 0.93 0.87 0.9 0.97 0.83
RF-ID-$R^{2}$ 0.93 0.93 0.93 0.9 0.73
RF-Bias-RMSE 0.93 0.57 0.9 0.9 0.8
RF-Bias-$R^{2}$ 0.83 0.63 0.93 0.87 0.8
RF-GroupBias-RMSE 0.87 0.85 0.91 0.90 0.86
RF-GroupBias-$R^{2}$ 0.89 0.86 0.91 0.88 0.86
LGBM-IID-RMSE 0.87 0.93 0.97 1.0 0.93
LGBM-IID-$R^{2}$ 1.0 0.97 0.97 0.93 0.97
LGBM-ID-RMSE 0.93 0.9 0.9 0.97 0.83
LGBM-ID-$R^{2}$ 0.9 0.93 0.83 0.97 0.93
LGBM-Bias-RMSE 0.9 0.7 0.8 0.9 0.77
LGBM-Bias-$R^{2}$ 0.93 0.7 0.77 0.9 0.83
LGBM-GroupBias-RMSE 0.89 0.85 0.90 0.88 0.85
LGBM-GroupBias-$R^{2}$ 0.89 0.86 0.97 0.88 0.87
XGBoost-IID-RMSE 1.0 0.93 0.93 1.0 0.87
XGBoost-IID-$R^{2}$ 0.93 0.93 0.93 0.97 0.93
XGBoost-ID-RMSE 0.93 0.9 1.0 0.93 0.97
XGBoost-ID-$R^{2}$ 0.93 0.9 0.93 0.93 0.87
XGBoost-Bias-RMSE 0.97 0.53 0.77 0.87 0.87
XGBoost-Bias-$R^{2}$ 0.83 0.5 0.9 0.87 0.83
XGBoost-GroupBias-RMSE 0.92 0.84 0.83 0.86 0.88
XGBoost-GroupBias-$R^{2}$ 0.87 0.82 0.90 0.84 0.88
CatBoost-IID-RMSE 0.87 0.97 0.97 1.0 1.0
CatBoost-IID-$R^{2}$ 0.87 0.97 0.93 0.93 1.0
CatBoost-ID-RMSE 0.93 0.87 0.93 0.97 0.97
CatBoost-ID-$R^{2}$ 0.97 0.9 0.93 0.97 1.0
CatBoost-Bias-RMSE 0.93 0.6 0.97 0.8 0.83
CatBoost-Bias-$R^{2}$ 0.93 0.63 0.93 0.8 0.87
CatBoost-GroupBias-RMSE 0.90 0.90 0.92 0.87 0.87
CatBoost-GroupBias-$R^{2}$ 0.90 0.87 0.90 0.86 0.88
Table 6: The ratio that the p-value of test is larger than or equal to 0.05 in the 30 repeated experiments for classification base learners. The name of each row is “BaseLearnerName”-“TestName”-“MetricName”. The name of each column is the scene name.
Ratio $p>=0.05$ withdraw higgs grid insurance climate alert
Linear-IID-ROCAUC 0.96 0.93 0.93 0.93 1.0 0.97
Linear-IID-ACC 0.93 0.97 0.9 0.9 0.97 0.97
Linear-ID-ROCAUC 1.0 0.97 0.87 0.9 1.0 0.97
Linear-ID-ACC 0.93 0.97 0.87 0.93 0.93 1.0
Linear-Bias-ROCAUC 0.77 0.83 0.83 0.9 0.9 0.9
Linear-Bias-ACC 0.83 0.87 0.87 0.8 0.83 0.93
Linear-GroupBias-ROCAUC 0.83 0.88 0.87 0.88 0.91 0.87
Linear-GroupBias-ACC 0.87 0.89 0.87 0.88 0.87 0.89
MLP-IID-ROCAUC 0.97 1.0 0.97 0.97 0.93 0.93
MLP-IID-ACC 0.93 0.93 0.97 0.9 0.97 0.97
MLP-ID-ROCAUC 0.8 0.87 0.77 0.87 0.87 0.7
MLP-ID-ACC 0.8 0.83 0.83 0.8 0.73 0.8
MLP-Bias-ROCAUC 0.43 0.57 0.53 0.53 0.63 0.47
MLP-Bias-ACC 0.67 0.6 0.57 0.57 0.43 0.6
MLP-GroupBias-ROCAUC 0.70 0.76 0.71 0.73 0.77 0.69
MLP-GroupBias-ACC 0.77 0.73 0.77 0.74 0.63 0.73
SVM-IID-ROCAUC 0.97 0.97 0.93 0.97 0.97 1.0
SVM-IID-ACC 0.97 0.97 0.97 0.97 1.0 1.0
SVM-ID-ROCAUC 0.73 0.5 0.9 0.6 0.83 0.8
SVM-ID-ACC 0.9 0.67 0.37 0.97 0.97 0.97
SVM-Bias-ROCAUC 0.4 0.27 0.1 0.37 0.53 0.47
SVM-Bias-ACC 0.97 0.37 0.43 0.93 0.8 0.93
SVM-GroupBias-ROCAUC 0.63 0.48 0.36 0.52 0.69 0.66
SVM-GroupBias-ACC 0.90 0.56 0.62 0.90 0.85 0.86
RF-IID-ROCAUC 0.97 0.97 0.9 0.97 0.97 0.97
RF-IID-ACC 0.93 0.97 0.87 0.93 0.9 0.9
RF-ID-ROCAUC 0.97 0.87 0.9 0.9 0.83 0.97
RF-ID-ACC 1.0 0.87 0.9 0.97 1.0 0.87
RF-Bias-ROCAUC 0.73 0.9 0.77 0.87 0.93 0.87
RF-Bias-ACC 0.9 0.93 0.8 0.97 0.8 0.7
RF-GroupBias-ROCAUC 0.85 0.87 0.84 0.9 0.91 0.86
RF-GroupBias-ACC 0.87 0.89 0.85 0.89 0.88 0.85
LGBM-IID-ROCAUC 1.0 0.97 0.97 0.97 0.87 1.0
LGBM-IID-ACC 0.93 0.97 0.93 0.9 1.0 0.9
LGBM-ID-ROCAUC 1.0 0.87 0.93 0.93 0.87 0.93
LGBM-ID-ACC 0.97 0.83 0.97 1.0 0.97 0.97
LGBM-Bias-ROCAUC 0.7 0.93 0.83 0.73 0.97 0.9
LGBM-Bias-ACC 0.83 0.9 0.77 0.83 0.8 0.8
LGBM-GroupBias-ROCAUC 0.81 0.88 0.85 0.89 0.91 0.86
LGBM-GroupBias-ACC 0.88 0.88 0.84 0.90 0.86 0.85
XGBoost-IID-ROCAUC 1.0 0.97 1.0 0.8 0.93 1.0
XGBoost-IID-ACC 0.93 1.0 0.97 1.0 0.93 0.87
XGBoost-ID-ROCAUC 0.9 0.83 0.9 0.9 0.87 1.0
XGBoost-ID-ACC 0.97 0.87 0.93 0.97 0.93 0.97
XGBoost-Bias-ROCAUC 0.83 0.97 0.83 0.9 0.83 0.83
XGBoost-Bias-ACC 0.87 0.93 0.9 0.83 0.8 0.87
XGBoost-GroupBias-ROCAUC 0.88 0.89 0.90 0.90 0.87 0.89
XGBoost-GroupBias-ACC 0.88 0.89 0.88 0.88 0.89 0.88
CatBoost-IID-ROCAUC 0.93 0.8 0.9 0.97 0.93 0.9
CatBoost-IID-ACC 1.0 0.9 1.0 0.87 1.0 0.97
CatBoost-ID-ROCAUC 0.97 0.93 0.93 0.93 0.87 0.97
CatBoost-ID-ACC 0.97 0.93 0.83 0.97 0.97 0.97
CatBoost-Bias-ROCAUC 0.83 0.87 0.8 0.8 0.9 0.87
CatBoost-Bias-ACC 0.87 0.9 0.93 0.87 0.8 0.8
CatBoost-GroupBias-ROCAUC 0.86 0.89 0.88 0.86 0.89 0.87
CatBoost-GroupBias-ACC 0.85 0.89 0.89 0.90 0.86 0.86
We also visualization the QQ plot of the baseline evaluation model in figure 5 and the assumption assessment result of our evaluation model in table 7. From the visualization, the error distribution of baseline is close to normal distribution. Although the HetEM-Linear and HetEM-CatBoost does not pass our ID check and Bias check, the dramatically error reduction of our evaluation model is still believed as valuable while it can be not used to calculate the upper bound of the 2 evaluation model directly in this scene.
Ratio $p\geq 0.05$ A-Share Trade
Linear-IID 1.0
Linear-ID 0.0
Linear-Bias 0.0
CatBoost-IID 1.0
CatBoost-ID 0.0
CatBoost-Bias 0.0
Table 7: The ratio that p-value of test is larger than or equal to 0.05 in the 5 repeated experiments for A-share trade scene. The name of each row is “BaseLearnerName”-“TestName”-“MetricName”. The name of each column is the scene name.
Figure 5: QQ plot of evaluation error of Baseline (Last10DayRoI).
Appendix D Hyper-parameters of evaluation model
There are 7 kinds of base learners for the evaluation model learning in our experiments. We use the default hyper-parameters for the 7 kinds of base learners in 6 scenes where evaluation metrics are RMSE and $R^{2}$ as following.
  "Linear Regression": {
    "copy_X": true,
    "fit_intercept": true,
    "positive": false
  "MLP": {
    "activation": "relu",
    "alpha": 0.0001,
    "batch_size": "auto",
    "beta_1": 0.9,
    "beta_2": 0.999,
    "early_stopping": false,
    "epsilon": 1e-08,
    "hidden_layer_sizes": [
      100
    "learning_rate": "constant",
    "learning_rate_init": 0.001,
    "max_fun": 15000,
    "max_iter": 200,
    "momentum": 0.9,
    "n_iter_no_change": 10,
    "nesterovs_momentum": true,
    "power_t": 0.5,
    "shuffle": true,
    "solver": "adam",
    "tol": 0.0001,
    "validation_fraction": 0.1,
    "verbose": false,
    "warm_start": false
  "SVM": {
    "C": 1.0,
    "cache_size": 200,
    "coef0": 0.0,
    "degree": 3,
    "epsilon": 0.1,
    "gamma": "scale",
    "kernel": "rbf",
    "max_iter": -1,
    "shrinking": true,
    "tol": 0.001,
    "verbose": false
  "RF": {
    "bootstrap": true,
    "ccp_alpha": 0.0,
    "criterion": "squared_error",
    "max_features": 1.0,
    "min_impurity_decrease": 0.0,
    "min_samples_leaf": 1,
    "min_samples_split": 2,
    "min_weight_fraction_leaf": 0.0,
    "n_estimators": 100,
    "verbose": 0,
    "warm_start": false
  "XGBoost": {
    "objective": "reg:squarederror",
    "missing": "nan",
    "n_estimators": 100
  "LightGBM": {
    "boosting_type": "gbdt",
    "colsample_bytree": 1.0,
    "importance_type": "split",
    "learning_rate": 0.1,
    "max_depth": -1,
    "min_child_samples": 20,
    "min_child_weight": 0.001,
    "min_split_gain": 0.0,
    "n_estimators": 100,
    "num_leaves": 31,
    "reg_alpha": 0.0,
    "reg_lambda": 0.0,
    "subsample": 1.0,
    "subsample_for_bin": 200000,
    "subsample_freq": 0
  "CatBoost": {
    "loss_function": "RMSE"
The default hyper-parameters of the 7 kinds of base learners in 5 scenes where evaluation metrics are ROC-AUC and ACC as following.
  "LogisticRegression": {
    "C": 1.0,
    "dual": false,
    "fit_intercept": true,
    "intercept_scaling": 1,
    "max_iter": 100,
    "multi_class": "deprecated",
    "penalty": "l2",
    "solver": "lbfgs",
    "tol": 0.0001,
    "verbose": 0,
    "warm_start": false
  "MLP": {
    "activation": "relu",
    "alpha": 0.0001,
    "batch_size": "auto",
    "beta_1": 0.9,
    "beta_2": 0.999,
    "early_stopping": false,
    "epsilon": 1e-08,
    "hidden_layer_sizes": [100],
    "learning_rate": "constant",
    "learning_rate_init": 0.001,
    "max_fun": 15000,
    "max_iter": 200,
    "momentum": 0.9,
    "n_iter_no_change": 10,
    "nesterovs_momentum": true,
    "power_t": 0.5,
    "shuffle": true,
    "solver": "adam",
    "tol": 0.0001,
    "validation_fraction": 0.1,
    "verbose": false,
    "warm_start": false
  "SVM": {
    "C": 1.0,
    "break_ties": false,
    "cache_size": 200,
    "coef0": 0.0,
    "decision_function_shape": "ovr",
    "degree": 3,
    "gamma": "scale",
    "kernel": "rbf",
    "max_iter": -1,
    "probability": false,
    "shrinking": true,
    "tol": 0.001,
    "verbose": false
  "RandomForest": {
    "bootstrap": true,
    "ccp_alpha": 0.0,
    "criterion": "gini",
    "max_features": "sqrt",
    "min_impurity_decrease": 0.0,
    "min_samples_leaf": 1,
    "min_samples_split": 2,
    "min_weight_fraction_leaf": 0.0,
    "n_estimators": 100,
    "oob_score": false,
    "verbose": 0,
    "warm_start": false
  "XGBoost": {
    "objective": "binary:logistic",
    "enable_categorical": false,
    "missing": "NaN"
  "LightGBM": {
    "boosting_type": "gbdt",
    "colsample_bytree": 1.0,
    "importance_type": "split",
    "learning_rate": 0.1,
    "max_depth": -1,
    "min_child_samples": 20,
    "min_child_weight": 0.001,
    "min_split_gain": 0.0,
    "n_estimators": 100,
    "num_leaves": 31,
    "reg_alpha": 0.0,
    "reg_lambda": 0.0,
    "subsample": 1.0,
    "subsample_for_bin": 200000,
    "subsample_freq": 0
  "CatBoost": {}
We also use the Grid search for the hyperparameter searching in quantum trade scene, the searched range is listed as following, the cross validation for the searching is 3-fold cross validation.
    "CatBoost" = {
        ‘iterations’: [100, 200],
        ‘depth’: [6, 8],
        ‘learning_rate’: [0.01, 0.1],
        ‘l2_leaf_reg’: [1, 3],
        ‘border_count’: [32, 64],
Appendix E Original evaluation errors
We listed the original average root mean square evaluation errors for 4 different evaluation metrics: ROC-AUC, ACC, RMSE, $R^{2}$ in the 11 scenarios for evaluation model learning.
Table 8: Evaluation error when evaluation metric is RMSE.
RMSE TERECO Calculus kin8nm nuscale ad
holdout-100 0.45471 0.26197 0.00499 0.01521 0.10611
holdout-50 0.28538 0.32978 0.00435 0.01328 0.19199
holdout-20 0.49894 0.29417 0.03387 0.01584 0.21959
holdout-10 0.39836 0.22317 0.00591 0.01706 0.21805
CV-5 0.43369 0.25795 0.00525 0.01518 0.07848
CV-10 0.39857 0.25568 0.00551 0.01513 0.06187
Bootstrap 0.46825 0.26353 0.00977 0.01544 0.13219
Linear 0.00474 0.00918 0.00027 0.00077 0.00555
MLP 0.03528 0.31249 0.02601 0.05205 0.03862
SVR 0.03881 0.04573 0.03729 0.02167 0.03665
RF 0.01348 0.00513 0.00193 0.00335 0.01378
LGBM 0.00998 0.00530 0.00219 0.00301 0.00985
XGBoost 0.01223 0.00549 0.00235 0.00331 0.01183
CatBoost 0.00756 0.00468 0.00159 0.00241 0.00695
Table 9: Evaluation error when evaluation metric is $R^{2}$.
RMSE TERECO Calculus kin8nm nuscale ad
holdout-100 0.12591 0.38410 0.01355 0.00946 0.19670
holdout-50 0.11428 0.53892 0.04429 0.01026 7.52594
holdout-20 0.20604 0.74010 0.03152 0.01594 6.24597
holdout-10 0.15633 0.70377 0.04148 0.02123 5.58360
CV-5 0.19360 0.42477 0.01200 0.00950 2.98845
CV-10 0.25662 0.43089 0.01052 0.00961 3.72466
Bootstrap 0.27573 0.39105 0.01911 0.01065 0.24769
Linear 0.01689 0.00061 0.00061 0.00174 0.01208
MLP 0.04252 0.02690 0.02690 0.05219 0.04072
SVR 0.06077 0.04628 0.04628 0.03578 0.05020
RF 0.04748 0.00441 0.00441 0.00770 0.03020
LGBM 0.03539 0.00508 0.00508 0.00702 0.02145
XGBoost 0.04258 0.00521 0.00521 0.00756 0.02537
CatBoost 0.02669 0.00368 0.00368 0.00553 0.01515
Table 10: Evaluation error when evaluation metric is ROC-AUC.
RMSE withdraw higgs grid insurance climate alert
holdout-100 0.02443 0.02132 0.01162 0.04982 0.06820 0.02185
holdout-50 0.03003 0.02323 0.01125 0.08180 0.07170 0.02442
holdout-20 0.03636 0.02878 0.01143 0.08537 0.07966 0.02515
holdout-10 0.04218 0.04217 0.01355 0.09063 0.12213 0.03166
CV-5 0.02445 0.02115 0.01161 0.06179 0.06894 0.02144
CV-10 0.02435 0.02118 0.01164 0.06180 0.06924 0.02159
Bootstrap 0.02634 0.02335 0.01236 0.04996 0.07509 0.02387
Linear 0.01587 0.01438 0.00723 0.01148 0.05123 0.01391
MLP 0.04397 0.04000 0.03974 0.05053 0.06654 0.04581
SVR 0.03433 0.02250 0.05526 0.03135 0.06141 0.07091
RF 0.02107 0.01608 0.01014 0.01645 0.05892 0.02255
LGBM 0.01934 0.01517 0.01056 0.01407 0.05555 0.01890
XGBoost 0.02116 0.01635 0.01016 0.01595 0.05970 0.02229
CatBoost 0.01748 0.01432 0.00888 0.01262 0.05232 0.01839
Table 11: Evaluation error when evaluation metric is ACC.
RMSE withdraw higgs grid insurance climate alert
holdout-100 0.05444 0.03244 0.01145 0.17974 0.10162 0.02764
holdout-50 0.06632 0.02550 0.01119 0.22168 0.13740 0.03271
holdout-20 0.06731 0.03507 0.01056 0.26858 0.16551 0.03930
holdout-10 0.06710 0.04374 0.01312 0.28142 0.19290 0.03623
CV-5 0.05443 0.03244 0.01145 0.17974 0.10145 0.02764
CV-10 0.05443 0.03244 0.01145 0.17973 0.10135 0.02765
Bootstrap 0.05489 0.03481 0.01266 0.17976 0.10362 0.02825
Linear 0.01850 0.01438 0.00696 0.01887 0.03357 0.01391
MLP 0.04886 0.03994 0.03924 0.05166 0.06719 0.04581
SVR 0.07261 0.03029 0.03749 0.07211 0.06634 0.07091
RF 0.03444 0.01641 0.00973 0.03068 0.05216 0.02255
LGBM 0.02929 0.01526 0.00972 0.02610 0.04359 0.02255
XGBoost 0.03395 0.01656 0.01000 0.02981 0.05219 0.01890
CatBoost 0.02655 0.01443 0.00879 0.02426 0.04645 0.01839
Appendix F Evaluation cost
We list the table to compare the average evaluation time of experimental (simulation-based) evaluation and our computation evaluation as following table 12.
Scenes Experiment (Simulation) Time Computation Time
Alert (Wilson et al. [2021]) 1 year, 8 months, and 15 days 0.176s
Withdraw (Wilson et al. [2023]) 1 year, 3 months, and 13 days 0.077s
Grid-S (Schäfer et al. [2016]) 1500s 0.198s
Higgs-S (Baldi et al. [2014]) (expected) 15 hours 0.151s
Insurance (Kumar [2023]) >1 year 0.500s
Climate-S (Lucas et al. [2013]) unknown 0.132s
TERECO (Li et al. [2022]) 34 weeks 0.014s
Calculus (Kramer et al. [2023]) 3 semesters 0.035s
Ad (Gokagglers [2018]) 3 campaigns 0.034s
kin8nm-S (Fischer [2022]) unknown 0.038s
NuScale-S (Huu Tiep [2024]) 1600 hours 0.046s
A-Share Trade 10 days 0.103s
Table 12: Expected evaluation time comparison in 12 scenes per subject.“S” means simulation.
Appendix G Upper bounds tables
We list the upper bound of generalized evaluation error and generalized causal effect evaluation error in the table 13 and table 14 for researchers’ quick query before data collection and upper bound estimation.
Samples/Confidence 0.5 0.8 0.95 0.99 0.999
10 $E+0.186$ $E+0.284$ $E+0.387$ $E+0.480$ $E+0.588$
20 $E+0.132$ $E+0.201$ $E+0.274$ $E+0.339$ $E+0.416$
30 $E+0.107$ $E+0.164$ $E+0.223$ $E+0.277$ $E+0.339$
100 $E+0.059$ $E+0.090$ $E+0.122$ $E+0.152$ $E+0.186$
1000 $E+0.0186$ $E+0.0284$ $E+0.0387$ $E+0.0480$ $E+0.0588$
10000 $E+0.00589$ $E+0.00897$ $E+0.0122$ $E+0.0151$ $E+0.0186$
100000 $E+0.00186$ $E+0.00284$ $E+0.00387$ $E+0.00480$ $E+0.00588$
1000000 $E+0.000589$ $E+0.000897$ $E+0.00122$ $E+0.00152$ $E+0.00186$
10000000 $E+0.000186$ $E+0.000284$ $E+0.000387$ $E+0.000480$ $E+0.000588$
100000000 $E+0.0000589$ $E+0.0000897$ $E+0.000122$ $E+0.000152$ $E+0.000186$
1000000000 $E+0.0000186$ $E+0.0000284$ $E+0.0000387$ $E+0.0000480$ $E+0.0000588$
Table 13: Upper bound of normalized generalized error. E is the empirical prediction error.
Samples/Confidence 0.5 0.8 0.95 0.99 0.999
10 $0.3722E+$ 0.372 $0.5672E+$ 0.567 $0.7742E+$ 0.774 $0.9602E+$ 0.960 $1.182E+$ 1.18
20 $0.2632E+$ 0.263 $0.4012E+$ 0.401 $0.5472E+$ 0.547 $0.6782E+$ 0.678 $0.8312E+$ 0.831
30 $0.2152E+$ 0.215 $0.3282E+$ 0.328 $0.4472E+$ 0.447 $0.5542E+$ 0.554 $0.6792E+$ 0.679
100 $0.1182E+$ 0.118 $0.1792E+$ 0.179 $0.2452E+$ 0.245 $0.3032E+$ 0.303 $0.3722E+$ 0.372
1000 $0.03722E+$ 0.0372 $0.05672E+$ 0.0567 $0.07742E+$ 0.0774 $0.09602E+$ 0.0960 $0.1182E+$ 0.118
10000 $0.01182E+$ 0.0118 $0.01792E+$ 0.0179 $0.02452E+$ 0.0245 $0.03032E+$ 0.0303 $0.03722E+$ 0.0372
100000 $0.003722E+$ 0.00372 $0.005672E+$ 0.00567 $0.007742E+$ 0.00774 $0.009602E+$ 0.00960 $0.01182E+$ 0.0118
1000000 $0.001182E+$ 0.00118 $0.001792E+$ 0.00179 $0.002452E+$ 0.00245 $0.003032E+$ 0.00303 $0.003722E+$ 0.00372
10000000 $0.0003722E+$ 0.000372 $0.0005672E+$ 0.000567 $0.0007742E+$ 0.000774 $0.0009602E+$ 0.000960 $0.001182E+$ 0.00118
100000000 $0.0001182E+$ 0.000118 $0.0001792E+$ 0.000179 $0.0002452E+$ 0.000245 $0.0003032E+$ 0.000303 $0.0003722E+$ 0.000372
1000000000 $0.00003722E+$ 0.0000372 $0.00005672E+$ 0.0000567 $0.00007742E+$ 0.0000774 $0.00009602E+$ 0.0000960 $0.0001182E+$ 0.000118
Table 14: Upper bound of normalized generalized causal effect error. E is the empirical prediction error.
Appendix H Paradigm of computational evaluation
In order to help user to better understand the paradigm of computational evaluation, we plot the decision tree to help evaluatology researchers conduct their own research as shown in the figure 6.
Figure 6: The decision tree that how to use computational evaluation models in your own evaluation task. The GEE means generalized evaluation error and the GCEE means generalized causal evaluation error. EC is evaluation condition, S is evaluation subject, and M is evaluation metric. P(M(S)|EC) is the potential outcome of evaluation metric after deploying the subject given evaluation condition.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Section 1
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 7
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Section 3
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: [Yes]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will push the source code, preprocessed data, intermediate data, generated data to the GitHub. And we will provide the link to download the aw data.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Appendix B, D
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Section 6
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix B
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [N/A]
  44. Justification: [Yes]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: It is mainly a theoretical research and it will not increase the gated release of models significantly because the computation requirement is not very high.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Despite the open data we used, we also scraped a dataset of A-share stocks in China by the AkShare with MIT License.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: [Yes]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: [Yes]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.